One gram of radium disintegrates to 0.2g in 3720 year. What is it`s half life period

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From the first order Rate equation we know,
log[A] – log[A0] = -kt
Putting the values we get,
log(1/226) – log(0.2/226) = -k*3720
k = 0.698/3720......................(1)
From the half life period:
t1/2 = 0.693/k
substituting from the equation (1) we get,
t1/2 = 0.693*3720/0.693 = 3720 years itself.